# The Latency Price of Threshold Cryptosystem in Blockchains

### v1

@@ Abstract
Threshold cryptography is essential for many blockchain protocols. For example, many protocols rely on threshold common coin to implement asynchronous consensus, leader elections, and provide support for randomized applications. Similarly, threshold signature schemes are frequently used for protocol efficiency and state certification, and threshold decryption and threshold time-lock puzzles are often necessary for privacy.
In this paper, we study the interplay between threshold cryptography and a class of blockchains that use Byzantine-fault tolerant (BFT) consensus protocols with a focus on latency. More specifically, we focus on blockchain-native threshold cryptosystem, where the blockchain validators seek to run a threshold cryptographic protocol once for every block with the block contents as an input to the threshold cryptographic protocol. All existing approaches for blockchain-native threshold cryptosystems introduce a latency overhead of at least one message delay for running the threshold cryptographic protocol. In this paper, we first propose a mechanism to eliminate this overhead for blockchain-native threshold cryptosystems with tight thresholds, i.e., in threshold cryptographic protocols where the secrecy and reconstruction thresholds are the same. However, many real-world proof-of-stake-based blockchain-native threshold cryptosystems rely on ramp thresholds, where reconstruction thresholds are strictly greater than secrecy thresholds. For these blockchains, we formally demonstrate that the additional delay is unavoidable. We then introduce a mechanism to minimize this delay in the optimistic case. We implement our optimistic protocol for the proof-of-stake distributed randomness scheme on the Aptos blockchain. Our measurements from the Aptos mainnet show that the optimistic approach reduces latency overhead by 71%, from 85.5 ms to 24.7 ms, compared to the existing method.
@@ 1 Introduction
Threshold cryptography plays a vital role in modern blockchains, where various applications rely on threshold cryptography primitives such as distributed randomness, threshold signature, and threshold decryption. In threshold cryptography, a secret is shared among a set of parties using a threshold secret sharing [41, 17], and parties seek to collaboratively evaluate a function of the shared secret and some public information without revealing the shared secret. For security, we require that the function of the shared secret and the public information is revealed only if a threshold fraction of parties contribute to the function evaluation.
@@
In this paper, we study the interplay between threshold cryptography and a class of blockchains that use Byzantine-fault tolerant (BFT) consensus protocols with a focus on latency. More specifically, we focus on blockchain-native threshold cryptosystem, where the blockchain validators seek to run a threshold cryptographic protocol $\sf TC$ once for every block with the block’s content as an input to $\sf TC$ protocol. In particular, we focus on schemes where the secret is shared using the Shamir secret sharing scheme [41], and the threshold cryptographic protocol is non-interactive, i.e., parties send a single message during the threshold cryptography protocol.
@@
One concrete example of a blockchain-native threshold cryptosystem is the recent distributed randomness protocol for proof-of-stake blockchains [23], which has been deployed in the Aptos blockchain [10]. In [23], parties collaboratively compute a threshold verifiable random function (VRF) to generate shared randomness for each block, using the cryptographic hash of the block as an input to the threshold VRF. Similarly, Kavousi et al. [33] propose to use threshold decryption to mitigate Maximal Extractable Value (MEV) attacks by the block proposers. Specifically, [33] presents a solution where blockchain validators first order a set of encrypted transactions using a consensus protocol. Next, upon ordering, block validators run a threshold decryption protocol to decrypt the finalized transactions and execute them. Another practical example is the use of threshold signatures (or multi-signatures) by several blockchains [10, 47] to certify the blockchain state after each block, and the aggregated threshold signature serves as proof of transaction inclusion for clients. In these blockchains, validators sign and exchange signature shares once each block is finalized and executed. This collective signing also prevents forking of the blockchain state in order-then-execute blockchains, particularly when execution is non-deterministic due to software bugs.
@@
One limitation of existing blockchain-native threshold cryptosystem is that parties participate in the $\sf TC$ protocol only after the block is finalized. Hence, all existing protocol introduces at least one additional message delay before the output of the $\sf TC$ protocol is available, for the parties to exchange their $\sf TC$ shares. As a result, blockchains that seek to use the output of the $\sf TC$ protocol to execute the finalized transactions also incur this additional latency. For blockchains with optimal consensus latency, the additional round of communication adds at least 33% latency overhead, which is significant. More precisely, the optimal consensus latency is known to be three message delays [38, 7, 36], which is achieved in several protocols and systems [21, 8]. For such protocols, the additional round of communication adds at least 33% latency overhead.
@@
In this paper, we investigate whether the additional delay is inherent in supporting threshold cryptography in BFT-based blockchains. More specifically, let $\sf TC$ be a threshold cryptography scheme. Then, the secrecy threshold of $\sf TC$ is the upper bound on the number of $\sf TC$ messages an adversary can learn without learning the output of the $\sf TC$ protocol. Alternatively, the reconstruction threshold is the number of $\sf TC$ messages an honest party requires to be able to compute the $\sf TC$ output. Committee-based blockchains where the parties have equal weights (stakes), such as Dfinity [31], can use blockchain-native threshold cryptosystem with the same secrecy threshold as the reconstruction threshold. For a wide variety of these blockchains, we present a simple protocol in which the parties can compute the $\sf TC$ output simultaneously with the block finalization time. More specifically, our protocol applies to all BFT consensus protocols that satisfy the property that a value is finalized if and only if a threshold number of parties prefinalize the value.
@@
However, many deployed proof-of-stake based blockchains [27, 43, 10, 47] where parties have unequal stakes, will rely on threshold cryptography with ramp thresholds [18], i.e., use threshold cryptographic protocols where the reconstruction threshold is strictly larger than the secrecy thresholds. The ramp nature of threshold cryptography in these protocols is because these protocols assign to each party an approximate number of shares proportional to their stake [48]. This approximate assignment of a number of shares to each party introduces a gap between the secrecy and reconstruction threshold, as the assignment process may allocate more shares to the corrupt parties and fewer shares to honest ones. Somewhat surprisingly, we prove a lower bound result illustrating that for blockchain-native threshold cryptosystem with ramp thresholds, the extra latency incurred by existing protocols is inherent for a wide family of consensus protocols.
@@
As a practical solution to circumvent this impossibility result, we propose a mechanism to design blockchain-native threshold cryptosystem protocols with ramp thresholds that achieve small latency overhead under optimistic executions. To demonstrate the effectiveness of our approach, we implement our optimistic solution atop the distributed randomness protocol (based on threshold VRF) used in the Aptos blockchain and evaluate its performance with their prior protocol. Our evaluation with real-world deployment illustrates that our optimistic approach reduces latency overhead by 71%.
To summarize, this paper makes the following contributions:
@@
  • We show an impossibility result (Theorem 6.1) indicating that the latency overhead is inherent for blockchain-native threshold cryptosystem with ramp thresholds, and present a solution (Algorithm 3) that can remove the latency overhead under optimistic scenarios. The results apply to proof-of-stake blockchain systems where parties have unequal weights.
  • We implement our solution of ramp thresholds for distributed randomness and present evaluation numbers from the Aptos mainnet deployment. The evaluation demonstrates that the solution significantly improves the latency overhead by $71\%$, from $85.5$ to $24.7$ ms.
2 Related Work
Threshold cryptography in blockchains. In many applications of modern blockchains, threshold cryptography plays a vital role. An important example of blockchain-native threshold cryptosystem is to generate distributed randomness for blockchains [23, 31], which samples a fresh randomness beacon for each finalized block, allowing any transaction in the block to directly sample randomness pseudo-randomly using the block-level beacon. For distributed randomness, existing real-world deployments [23, 31, 10, 47, 28] follow the threshold VRF-based approach. Certifying the blockchain state using threshold signatures (or multi-signature) is also a common practice [10, 47] for order-then-execute blockchains. Numerous blockchain research [40, 49, 13, 51, 39, 33, 37] focuses on MEV countermeasures and privacy enhancement using threshold decryption, where transactions are encrypted with threshold decryption and only revealed and executed by blockchain parties after finalization.
To the best of our knowledge, all existing constructions supporting the blockchain-native threshold cryptosystem mentioned above incur an additional round of latency, as parties participate in the threshold cryptographic protocol only after each block is finalized. We believe this is due to the black-box use of the underlying consensus mechanism. Looking ahead, our proposed constructions reduce this latency by using the underlying consensus protocols in a non-black-box manner. Our techniques are not tied to a specific blockchain but are applicable to a wide range of blockchains that satisfy Definition 5.
@@ 3 Preliminaries
Notations. For any integer $a$, we use $[a]$ to denote the ordered set $\{1,2,\ldots,a\}$. For any set $S$, we use $|S|$ to denote the size of set $S$. We use $\lambda$ to denote the security parameter. A machine is probabilistic polynomial time (PPT) if it is a probabilistic algorithm that runs in time polynomial in $\lambda$. We use ${\sf negl}(\lambda)$ to denote functions that are negligible in $\lambda$. We summarize the notations we use in the paper in Table 1.
@@
$\sf MBB_{FT}$ Multi-shot Byzantine Broadcast with finalization threshold (Definition 5)
@@
$\sf BTC_{FT}$ blockchain-native threshold cryptosystem with finalization threshold (Definition 13)
@@
Table 1: Table of Notations
3.1 System Model
We consider a set of $n$ parties labeled $1,2,...,n$, where each party executes as a state machine. For brevity, we present the results for parties with equal weights, which can be easily extended to the case with unequal weights. The parties communicate with each other by message passing, via pairwise connected communication channels that are authenticated and reliable. We consider a static adversary $\mathcal{A}$ that can corrupt up to $t$ parties before the execution of the system. A corrupted party can behave arbitrarily, and a non-corrupted party behaves according to its state machine. We say that a non-corrupted party is honest. We use $\mathcal{C}$ to denote the set of corrupted parties, and ${\mathcal{H}}$ to denote the set of honest parties. The network is assumed to be partially synchronous, where there exists a known message delay upper bound $\Delta$, and a global stabilization time (GST) after which all messages between honest parties are delivered within $\Delta$ [26]. The adversary can receive messages from any party instantaneously.
@@ 3.2 Blockchain Definitions
We define Multi-shot Byzantine Broadcast under partial synchrony as follows to capture the consensus layer of many real-world blockchains that assume a partial synchronous network. We will use Multi-shot Byzantine Broadcast and consensus interchangeably throughout the paper. Intuitively, Multi-shot Byzantine Broadcast consists of infinite instances of Byzantine Broadcast with rotating broadcasters and guarantees a total ordering among all instances.
@@
  • Termination. After GST, for any round $r\in\mathbb{N}$ each honest party eventually outputs ${\sf finalize}(r,m)$ where $m\in{\mathcal{M}}\cup\{\bot\}$.
  • Validity. If the designated broadcaster $B_{r}$ of round $r$ is honest and calls ${\sf bcast}(r,m)$ for any $m\in{\mathcal{M}}$ after GST, then all honest parties eventually output ${\sf finalize}(r,m)$.
@@
A Multi-shot Byzantine Broadcast protocol $\sf MBB$ defines the state machine for each party to solve Multi-shot Byzantine Broadcast.
The definition of Multi-shot Byzantine Broadcast captures many existing partially synchronous leader-based BFT protocols, or with minor modifications  [^1] Many chained BFT protocols such as HotStuff [50] and Jolteon [29] achieves a weaker Validity property. In these protocols, the finalization of the message proposed by the broadcaster of round $r$ requires multiple consecutive honest broadcasters starting from round $r$. This weaker Validity does not affect the results we present in this paper., such as PBFT [20], Tendermint [19], SBFT [30], HotStuff [50], Streamlet [22], Fast-Hotstuff [32], Jolteon [29] Moonshot [24] and many others. Multi-shot Byzantine Broadcast also captures another series of DAG-based BFT protocols, such as Bullshark [46], Shoal [45], Shoal++ [12], Cordial miners [34] and Mysticeti [14].
Now we define an execution of a $\sf MBB$ protocol, and when a message is globally finalized for a round in an execution.
Definition 2 (Execution)
@@
Definition 3 (Multivalent and univalent state)
During the execution of a Multi-shot Byzantine Broadcast protocol, the system is in a multivalent state for round $r$, if there exist two possible executions $\mathcal{E}\neq\mathcal{E}^{\prime}$ both extending the current configuration, where some honest party output differently in $\mathcal{E},\mathcal{E}^{\prime}$; the system is in a univalent state of $m\in{\mathcal{M}}\cup\{\bot\}$ for round $r$, if for all execution extending the current configuration, all honest party always outputs ${\sf finalize}(r,m)$.
@@
Intuitively, a message is globally finalized for a round $r$ in Multi-shot Byzantine Broadcast when it is the only message that can be the output of $r$. Global finalization is a global event that may not be immediately known to any honest party, but must occur no later than the moment that any honest party outputs ${\sf finalize}(r,\cdot)$. Compared to local finalization, which occurs when any honest party outputs ${\sf finalize}(r,\cdot)$, global finalization is more fundamental. A $\sf MBB$ protocol can easily alter the time of local finalization simply by redefining the conditions for an honest party to local finalize (such as adding artificial delays to local finalization). However, modifying the global finalization time requires a more comprehensive protocol change. Therefore, we use the global finalization time for our impossibility result. For the feasibility results where we present protocol constructions, the time of local finalization will be considered.
3.2.1 Multi-shot Byzantine Broadcast with Finalization Threshold.
The paper focuses on a family of $\sf MBB$ protocols that have a finalization threshold defined as follows. We call such a protocol Multi-shot Byzantine Broadcast with finalization threshold, or $\sf MBB_{FT}$.
Definition 5 (Finalization Threshold)
For any round $r\in\mathbb{N}$, the Multi-shot Byzantine Broadcast with finalization threshold, or $\sf MBB_{FT}$, has a state where a party calls ${\sf prefinalize}(r,m)$ to prefinalize a message $m\in{\mathcal{M}}\cup\{\bot\}$ for $r$, and sends a message $({\tt PREFIN},r,m)$ to all parties, such that
  • For any round $r$, any honest party calls ${\sf prefinalize}(r,m)$ for at most one $m\in{\mathcal{M}}$.
  • For any round $r$, any honest party can call ${\sf prefinalize}(r,\bot)$ after calling ${\sf prefinalize}(r,m)$ for some $m\in{\mathcal{M}}$, but not the reverse.
  • $m\in{\mathcal{M}}\cup\{\bot\}$ is globally finalized for $r$, if and only if there exist $t_{\sf fin}$ parties (or equivalently $t_{\sf fin}-|\mathcal{C}|$ honest parties) that have called ${\sf prefinalize}(r,m)$.
Any party outputs ${\sf finalize}(r,m)$ to locally finalize a message $m\in{\mathcal{M}}\cup\{\bot\}$ for a round $r$ when the party receives $({\tt PREFIN},r,m)$ messages from $t_{\sf fin}$ parties, which implies that $m$ is globally finalized for $r$.
@@
Intuitively, prefinalization is a local state of parties that satisfy the property that when enough honest parties prefinalize a message, then the message is globally finalized. A single party prefinalizing a message does not guarantee that the message will be finalized, and the party may finalize another message at the end. In many BFT protocols such as HotStuff [50], prefinalization is also named lock.
Examples of Multi-shot Byzantine Broadcast with finalization threshold. A larger number of $\sf MBB$ protocols used in partially synchronous blockchains fall into this family. As a concrete example, Jolteon [29] is a partially synchronous $\sf MBB_{FT}$ protocol deployed by blockchains such as Aptos [10] and Flow [28]. We explain in detail how Jolteon satisfies the definition of Multi-shot Byzantine Broadcast with finalization threshold with finalization threshold in Section 7.2. Other than Jolteon, numerous partially synchronous BFT protocols are also part of this family or can be easily adapted to fit into this family, such as PBFT [20], Tendermint [19], SBFT [30], HotStuff [50], Streamlet [22], Fast-Hotstuff [32], Moonshot [24] and many others. Additionally, another series of DAG-based consensus protocols also satisfy the definition of Multi-shot Byzantine Broadcast with finalization threshold with finalization threshold, such as Bullshark [46], Shoal [45], Shoal++ [12], Cordial miners [34] and Mysticeti [14].

### v3

@@ Abstract
Threshold cryptography is essential for many blockchain protocols. For example, many protocols rely on threshold common coin to implement asynchronous consensus, leader elections, and randomized applications. Similarly, threshold decryption and threshold time-lock puzzles are often necessary for privacy.
In this paper, we study the interplay between threshold cryptography and a class of blockchains that use Byzantine-fault tolerant (BFT) consensus protocols with a focus on latency. More specifically, we focus on blockchain-native threshold cryptosystem, where the blockchain validators seek to run a threshold cryptographic protocol once for every block with the block contents as an input to the threshold cryptographic protocol. All existing approaches for blockchain-native threshold cryptosystems introduce a latency overhead of at least one message delay for running the threshold cryptographic protocol. In this paper, we first propose a mechanism to eliminate this overhead for blockchain-native threshold cryptosystems with tight thresholds, i.e., in threshold cryptographic protocols where the secrecy and reconstruction thresholds are the same. However, real-world proof-of-stake-based blockchain-native threshold cryptosystems rely on ramp thresholds, where reconstruction thresholds are strictly greater than secrecy thresholds. For these blockchains, we demonstrate that the additional delay is unavoidable. We then introduce a mechanism to minimize this delay in the optimistic case. We implement our optimistic protocol for the proof-of-stake distributed randomness scheme on the Aptos blockchain. Our measurements from the Aptos mainnet show that the optimistic approach reduces latency overhead by 71%, from 85.5 ms to 24.7 ms, compared to the existing method.
@@ 1 Introduction
Threshold cryptography plays a vital role in modern blockchains, where various applications rely on primitives such as distributed randomness and threshold decryption. In threshold cryptography, a secret is shared among a set of parties using a threshold secret sharing [41, 18], and parties seek to collaboratively evaluate a function of the shared secret and some public input without revealing the shared secret. For security, the function of the shared secret and the public information is revealed only if a threshold fraction of parties contribute to the function evaluation.
@@
In this paper, we study the interplay between threshold cryptography and a class of blockchains that use Byzantine-fault tolerant (BFT) consensus protocols with a focus on latency. More specifically, we focus on blockchain-native threshold cryptosystem, where the blockchain validators seek to run a threshold cryptographic protocol $\sf TC$ once for every block with the block’s content as an input to $\sf TC$ protocol. We focus on schemes where the secret is shared using the Shamir secret sharing scheme [41], and the threshold cryptographic protocol is non-interactive, i.e., parties send a single message during the threshold cryptography protocol.
@@
One concrete example of a blockchain-native threshold cryptosystem is the recent distributed randomness protocol for proof-of-stake blockchains [23], that has been deployed in the Aptos blockchain [11]. In [23], parties collaboratively compute a threshold verifiable random function (VRF) to generate shared randomness for each block, using the cryptographic hash of the block as an input to the threshold VRF. Similarly, Kavousi et al. [33] propose to use threshold decryption to mitigate Maximal Extractable Value (MEV) attacks by the block proposers. Specifically, in [33] blockchain validators first order a set of encrypted transactions using a consensus protocol. Next, upon ordering, block validators run a threshold decryption protocol to decrypt the finalized transactions and execute them.
@@
One limitation of existing blockchain-native threshold cryptosystem is that parties participate in the $\sf TC$ protocol only after the block is finalized. Hence, all existing protocol introduces at least one additional message delay before the output of the $\sf TC$ protocol is available, for the parties to exchange their $\sf TC$ shares. As a result, blockchains that seek to use the output of the $\sf TC$ protocol to execute the finalized transactions also incur this additional latency. For blockchains [21, 8] with optimal consensus latency of three-message delay [38, 7, 36], the additional round of communication adds at least 33% latency overhead, which is significant.
@@
This paper studies whether the additional delay is inherent to support threshold cryptography in BFT-based blockchains. More specifically, let $\sf TC$ be a threshold cryptography scheme. Then, the secrecy threshold of $\sf TC$ is the upper bound on the number of $\sf TC$ messages an adversary can learn without learning the output of the $\sf TC$ protocol. Alternatively, the reconstruction threshold is the number of $\sf TC$ messages an honest party requires to be able to compute the $\sf TC$ output. Committee-based blockchains where the parties have equal weights (stakes), such as Dfinity [31], can use blockchain-native threshold cryptosystem with the same secrecy threshold as the reconstruction threshold. For a wide variety of these blockchains, we present a protocol in which the parties can compute the $\sf TC$ output simultaneously with the block finalization time. More specifically, our protocol applies to all BFT consensus protocols in which a value is finalized if and only if a threshold number of parties prefinalize the value.
@@
However, many proof-of-stake blockchains [27, 44, 11, 48] where parties have unequal stakes, will rely on threshold cryptography with ramp thresholds [19], i.e., use threshold cryptographic protocols where the reconstruction threshold is strictly larger than the secrecy thresholds. The ramp nature of threshold cryptography in these protocols is because these protocols assign to each party an approximate number of shares proportional to their stake [49]. This approximate assignment of a number of shares to each party introduces a gap between the secrecy and reconstruction threshold, as the assignment process may allocate more shares to the corrupt parties and fewer shares to honest ones. Somewhat surprisingly, we prove a lower bound result illustrating that for blockchain-native threshold cryptosystem with ramp thresholds, the extra latency incurred by existing protocols is inherent for a wide family of consensus protocols.
@@
To circumvent this impossibility result, we propose a mechanism to design blockchain-native threshold cryptosystem protocols with ramp thresholds that achieve small latency overhead under optimistic executions. To demonstrate the effectiveness, we implement our solution atop the distributed randomness protocol (based on threshold VRF) used in the Aptos blockchain and evaluate its performance with their prior protocol. Our evaluation with real-world deployment illustrates that our optimistic approach reduces latency overhead by 71%.
In summary, we make the following contributions:
@@
  • We prove an impossibility result (Theorem 5.1) indicating that the latency overhead is inherent for blockchain-native threshold cryptosystem with ramp thresholds, and present a solution (Algorithm 3) that can remove the latency overhead under optimistic scenarios. The results apply to proof-of-stake blockchain systems where parties have unequal weights.
  • We implement our solution of ramp thresholds for distributed randomness and present evaluation numbers from the Aptos mainnet deployment. The evaluation demonstrates that the solution significantly improves the randomness generation latency overhead by $71\%$, from $85.5$ to $24.7$ ms.
@@ 2 Preliminaries
Notations. For any integer $a$, we use $[a]$ to denote the ordered set $\{1,2,\ldots,a\}$. For any set $S$, we use $|S|$ to denote the size of set $S$. We use $\lambda$ to denote the security parameter. A machine is probabilistic polynomial time (PPT) if it is a probabilistic algorithm that runs in time polynomial in $\lambda$. We use ${\sf negl}(\lambda)$ to denote functions that are negligible in $\lambda$. We summarize the notations in Table 1.
@@
$\sf MBB_{FT}$ $\sf MBB$ with finalization threshold (Definition 4)
@@
$\sf BTC_{FT}$ $\sf BTC$ with finalization threshold (Definition 10)
@@
Table 1: Table of Notations.
System Model. We consider a set of $n$ parties labeled $1,2,...,n$, where each party executes as a state machine. For brevity, we present the results for parties with equal weights, which can be easily extended to the case with unequal weights. The parties communicate with each other by message passing, via pairwise connected communication channels that are authenticated and reliable. We consider a static adversary $\mathcal{A}$ that can corrupt up to $t$ parties before the execution of the system. A corrupted party can behave arbitrarily, and a non-corrupted party behaves according to its state machine. We say that a non-corrupted party is honest. We use $\mathcal{C}$ to denote the set of corrupted parties, and ${\mathcal{H}}$ to denote the set of honest parties. The network is assumed to be partially synchronous, where there exists a known message delay upper bound $\Delta$, and a global stabilization time (GST) after which all messages between honest parties are delivered within $\Delta$ [26]. The adversary can receive messages from any party instantaneously.
@@ 2.1 Blockchain Definitions
We define Multi-shot Byzantine Broadcast under partial synchrony as follows to capture the consensus layer of many real-world blockchains that assume a partial synchronous network. We will use Multi-shot Byzantine Broadcast and consensus interchangeably throughout the paper. Intuitively, Multi-shot Byzantine Broadcast consists of infinite instances of Byzantine Broadcast with rotating broadcasters and guarantees a total ordering among all instances. The primary reason for introducing a new definition, rather than relying on existing ones such as Byzantine Atomic Broadcast, is the necessity of incorporating rounds into the definition, as rounds will be referenced in later definitions (Definition 4, Definition 8). The definition of Multi-shot Byzantine Broadcast captures many existing partially synchronous leader-based BFT protocols, or with minor modifications  [^1] Many chained BFT protocols such as HotStuff [51] and Jolteon [29] achieve a weaker Validity property. In these protocols, the finalization of the message proposed by the broadcaster of round $r$ requires multiple consecutive honest broadcasters starting from round $r$. This weaker Validity does not affect the results we present in this paper., such as [21, 20, 30, 51, 22, 32, 29, 24], as well as DAG-based BFT protocols, such as [47, 46, 12, 34, 14].
@@
  • Termination. After GST, for any round $r\in\mathbb{N}$ each honest party $i$ eventually outputs ${\sf finalize}(r,m_{i})$ where $m_{i}\in{\mathcal{M}}\cup\{\bot\}$.
  • Validity. If the broadcaster $B_{r}$ of round $r$ is honest and calls ${\sf bcast}(r,m)$ for any $m\in{\mathcal{M}}$ after GST, then all honest parties eventually output ${\sf finalize}(r,m)$.
@@
A Multi-shot Byzantine Broadcast protocol $\sf MBB$ defines the state machine for each party to solve Multi-shot Byzantine Broadcast. Now we define an execution of an $\sf MBB$ protocol, and the globally finalization of a message for a round.
Definition 2 (Execution, multivalent and univalent state)
@@
During the execution of a Multi-shot Byzantine Broadcast protocol, the system is in a multivalent state for round $r$, if there exist two possible executions $\mathcal{E}\neq\mathcal{E}^{\prime}$ both extending the current configuration, where some honest party output differently in $\mathcal{E},\mathcal{E}^{\prime}$; the system is in a univalent state of $m\in{\mathcal{M}}\cup\{\bot\}$ for round $r$, if for all executions extending the current configuration, all honest parties always outputs ${\sf finalize}(r,m)$.
@@
Intuitively, a message is globally finalized for a round $r$ in Multi-shot Byzantine Broadcast when it is the only message that can be the output of $r$. Global finalization is a global event that may not be immediately known to any honest party, but must occur no later than the moment that any honest party outputs ${\sf finalize}(r,\cdot)$. Compared to local finalization, which occurs when any honest party outputs ${\sf finalize}(r,\cdot)$, global finalization is more fundamental.
Multi-shot Byzantine Broadcast with Finalization Threshold. The paper focuses on a family of $\sf MBB$ protocols that have a finalization threshold $t_{\sf fin}$ defined as follows, where $t_{\sf fin}$ is a parameter of the definition. We call such a protocol Multi-shot Byzantine Broadcast with finalization threshold, or $\sf MBB_{FT}$.
Definition 4 (Finalization Threshold $t_{\sf fin}$)
For any round $r\in\mathbb{N}$, the Multi-shot Byzantine Broadcast with finalization threshold, or $\sf MBB_{FT}$, has a step where a party prefinalizes $m$ for round $r$ by sending $({\tt PREFIN},r,m)$ to all parties, such that
  • For any round $r$, there exists at most one $m\in{\mathcal{M}}$ such that any honest party may send $({\tt PREFIN},r,m)$.
  • For any round $r$, any honest party can send $({\tt PREFIN},r,\bot)$ after sending $({\tt PREFIN},r,m)$ for some $m\in{\mathcal{M}}$, but not the reverse.
  • $m\in{\mathcal{M}}\cup\{\bot\}$ is globally finalized for $r$, if and only if there exist $t_{\sf fin}$ parties (or equivalently $t_{\sf fin}-|\mathcal{C}|$ honest parties) that have sent $({\tt PREFIN},r,m)$.
Any honest party outputs ${\sf finalize}(r,m)$ to locally finalize a message $m\in{\mathcal{M}}\cup\{\bot\}$ for a round $r$ when the party receives $({\tt PREFIN},r,m)$ messages from $t_{\sf fin}$ parties, which implies that $m$ is globally finalized for $r$.
@@
Intuitively, prefinalization is a local state of any party. When enough honest parties prefinalize a message, the message is globally finalized. A single party prefinalizing a message does not guarantee that the message will be finalized, and the party may finalize another message at the end. In many BFT protocols such as HotStuff [51] and Jolteon [29], prefinalization is also named lock. In PBFT [21], party prefinalizes a message when sending a commit for the message.
Examples of Multi-shot Byzantine Broadcast with finalization threshold.
A larger number of $\sf MBB$ protocols used in partially synchronous blockchains fall into this family. As a concrete example, Jolteon [29] is a partially synchronous $\sf MBB_{FT}$ protocol deployed by blockchains such as Aptos [11] and Flow [28]. We explain in detail how Jolteon satisfies the definition of Multi-shot Byzantine Broadcast with finalization threshold with finalization threshold in Section 0.E.2. Other than Jolteon, numerous partially synchronous BFT protocols are also part of this family or can be easily adapted to fit into this family, such as PBFT [21], Tendermint [20], SBFT [30], HotStuff [51], Fast-Hotstuff [32], Moonshot [24] and many others. Additionally, another series of DAG-based consensus protocols also satisfy the definition of Multi-shot Byzantine Broadcast with finalization threshold with finalization threshold, such as Shoal++ [12], Sailfish [43], Cordial miners [34] and Mysticeti [14].